A 10.0 x 10^(-6) F capacitor is charged by a 10.0V battery through a resistance R. The capacitor reaches a potential difference of 4.00V in a time interval of 3.00 s after charging begins. Find R
1
Expert's answer
2021-02-16T10:32:15-0500

Capacitance, "\text{C}= 10\times 10^{-6}\text{F}=10^{-5}\text{F}"

Voltage, "\text{V}=10 \text{volts}"


As the Capacitor reaches 4 volts in 3 sec

Then the charge stored in the capacitorin 3 sec is

"\text{Q=CV}\\Q=10^{-5}\times 4=4\times 10^{-5}\text{Columb}"


Total charge on the capacitor, "Q_o=CV=10^{-5}\times 10=10^{-4}\text{columb}"


Using the capacitor Charging equation,

"Q=Q_oe^{-\frac{t}{\tau}}"

"\Rightarrow 10^{-5}=10^{-4}e^{-\frac{3}{\tau}}"

"\Rightarrow 0.1=e^{-\frac{3}{\tau}}"

Taking log on Both sides and solving-


"\Rightarrow \dfrac{-3}{\tau}=log0.1"


"\Rightarrow \tau=\dfrac{-3}{-1}=3"


As we know Time constant, "\tau=RC"

"\Rightarrow R=\dfrac{\tau}{C}=\dfrac{3}{10^{-5}}=3\times 10^5 \Omega"